find a quadratic equation having 5/3 and 1/3 as roots
Answer by ankor@dixie-net.com(22740)

find a quadratic equation having 5/3 and 1/3 as roots
Derive factors from the roots
x = 5/3
(3x - 5) = 0
and
x = 1/3
(3x - 1) = 0
FOIL
(3x - 5)(3x - 1) = 9x^2 - 3x - 15x + 5 = 3x^2 - 18x + 5: is the equation